There are 850 Douglas fir and ponderosa pine trees in a section of forest bought by Sawz Logging Co. The company paid an average of $300 for each Douglas fir and $225 for each ponderosa pine. If the company paid $217,500 for the trees, how many of each kind did the company buy?
How many Douglas firs?


How many Ponderosa Pine?

Answer by checkley71(8403)

D+P=850 OR D=850-P
300D+225P=217500 NOW REPLACE D BY (850-P)
300(850-P)+225P=217500
255000-300P+225P=217500
-75P=217500-255000
-75P=-37500
P=-37500/-75
P=500 NUMBER OF PONDEROSA PINETREES SOLD.
D+500=850
D=850-500
D=350 NUMBER OF DOUGLAS TREES SOLD.
PROOF
300*350+500*225=217500
105000+112500=217500
217500=217500